OpenStudy (anonymous):

Determine the number and type of solutions for 5x2 + 1x + 7 = 0. 1 Rational Double 2 Irrational 2 Rational 2 Complex

OpenStudy (ipwnbunnies):

You'll have to use that discriminant I gave ya. And this link is helpful: http://www.mathwarehouse.com/quadratic/discriminant-in-quadratic-equation.php Particularly the first half.

OpenStudy (anonymous):

5x^2 + x + 7 = 0 The Discriminant D = b^2 - 4ac = 1 - 140 = - 139 < 0 There are 2 complex roots.